OpenStudy (samigupta8):

Helium gas taken from a nuclear experiment contains the isotopes 3He and 4He.One mole of the gas weighs 3.25 g and has a pressure of 1 atm.What is the partial pressure of 3He?

OpenStudy (samigupta8):

@priyar

OpenStudy (priyar):

Ptotal = XaPa +XbPb where a- 3He and b-4He

OpenStudy (samigupta8):

Bt we don't know the mole fraction

OpenStudy (samigupta8):

And even the partial pressure of other isotope is also not given

OpenStudy (priyar):

use this to find x: (x->wt. of a ) \[\frac{ x }{ 3 } + \frac{ 3.25-x }{ 4 } =1 \] after u get x you can easily find the mole fraction of each.. that is Xa and Xb.. And you are already given Ptotal..

OpenStudy (priyar):

are you sure Pa is not given?

OpenStudy (samigupta8):

Yep !

OpenStudy (kainui):

I don't see how you can solve this without more information, since I'm looking at for 3 and 4 referring to the He isotopes respectively, \[1= X_3+X_4\]\[P_{tot} = P_3X_3 + P_4X_4 \]\[M_{tot} = M_3X_3 + M_4X_4 \] Essentially you have 3 equations and 4 unknowns, so we need one more bit of information. Is there any other info, can you make sure you have copied the question down exactly as it is written and have not left out any other information?

OpenStudy (samigupta8):

Sir ! I have checked it but they didn't give more information of the question than the posted one...

OpenStudy (samigupta8):

And btw...the formula u all are giving is Raoult's law for mixing of two ideal solution

OpenStudy (kainui):

Is Raoult's law bad?

OpenStudy (samigupta8):

Think so ! Bcoz after all it is used for ideal solutions of liquids and not gases...